# HEPPO: Hardware-Efficient Proximal Policy Optimization A Universal Pipelined Architecture for Generalized Advantage Estimation

### v1

@@ Abstract
This paper introduces HEPPO, an FPGA-based accelerator designed to optimize the Generalized Advantage Estimation (GAE) stage in Proximal Policy Optimization (PPO). Unlike previous approaches that focused on trajectory collection and actor-critic updates, HEPPO addresses GAE’s computational demands with a parallel, pipelined architecture implemented on a single System-on-Chip (SoC). This design allows for the adaptation of various hardware accelerators tailored for different PPO phases. A key innovation is our strategic standardization technique, which combines dynamic reward standardization and block standardization for values, followed by 8-bit uniform quantization. This method stabilizes learning, enhances performance, and manages memory bottlenecks, achieving a 4x reduction in memory usage and a 1.5x increase in cumulative rewards. We propose a solution on a single SoC device with programmable logic and embedded processors, delivering throughput orders of magnitude higher than traditional CPU-GPU systems. Our single-chip solution minimizes communication latency and throughput bottlenecks, significantly boosting PPO training efficiency. Experimental results show a 30% increase in PPO speed and a substantial reduction in memory access time, underscoring HEPPO’s potential for broad applicability in hardware-efficient reinforcement learning algorithms.
@@ I-C Related Work
This section discusses key contributions in hardware acceleration for RL, aligning with our work on HEPPO by addressing the computational demands of RL algorithms through innovative hardware designs, quantization techniques, and memory management strategies.
@@ II-C2 Quantization of Values
III HEPPO Architecture Details
@@ III-C System Architectue
Figure 5 (a) shows the micro-architecture of HEPPO, which consists of Rewards Loaders (ReLs), Values Loaders (VaLs), and compute Processing Elements (PEs) forming a one-dimensional systolic array with N rows.
Figure 5: The HEPPO architecture consisting of Rewards Loaders (ReLs), Values Loaders (VaLs), PEs, system crossbar, and BRAM stack memory
Parallelization: Rows in the systolic array run concurrently and independently, each processing distinct vectors from different agents assigned by a round-robin fashion. When one row finishes, it gets a new set of vectors. This parallel architecture enhances HEPPO’s efficiency and scalability. While the BRAM stack memory enables substantial data transfers, a crossbar network ensures robust connections between ReLs, VaLs, and PEs to the BRAM stack memory.
@@ V-C Summary of Rewards and Values Quantization Approaches
V-D Hardware Implementation of HEPPO
A parameterized Verilog model of HEPPO’s proposed pipelined architecture, as described in Subsection III-C, has been developed with a data width of 32 bits (after fetching and de-quantizing the elements). The AMD-Xilinx Zynq® UltraScale+™ MPSoC ZCU106 Evaluation Kit was chosen to host our implementation. This device integrates a quad-core Arm® Cortex™-A53 processing system (PS) and a dual-core Arm Cortex-R5 real-time processor, providing the necessary computing for running the environment. The FPGA fabric within the Programmable Logic (PL) provides extensive resources for custom logic implementation, including neural network inference and GAE computation.
@@ VI Conclusion and Future Work
We introduced HEPPO (Hardware-Efficient Proximal Policy Optimization), an FPGA-based implementation designed to accelerate the GAE stage of the PPO algorithm. HEPPO utilizes dynamic reward standardization and 8-bit uniform quantization, reducing memory usage by 4x and increasing cumulative rewards by 50%.
@@
HEPPO leverages AMD-Xilinx Zynq UltraScale+ MPSoC’s capabilities, integrating environment simulation, neural network inference, backpropagation, and GAE computation on a single SoC. This minimizes communication latency and optimizes data handling which originally accounted for around 11% of the training.
Further incremental optimizations of HEPPO’s SoC are possible. Overclocking techniques [18] and bit-serial computation [19] in FPGAs can be employed to accelerate overall processing. To mitigate power consumption, multiple clock domains can be implemented for the ARM cores, the DNN, and the GAE calculations. High-performance clock-domain crossing (CDC) FIFOs can facilitate faster data transfers [20, 21, 22, 23, 24]. Additionally, hardware-efficient data compression methods, optimized for deep learning workloads, can be leveraged to minimize data transfers [25, 26, 27].

### v2

# HEPPO-GAE: Hardware-Efficient Proximal Policy Optimization with Generalized Advantage Estimation

@@ Abstract
This paper introduces HEPPO-GAE, an FPGA-based accelerator designed to optimize the Generalized Advantage Estimation (GAE) stage in Proximal Policy Optimization (PPO). Unlike previous approaches that focused on trajectory collection and actor-critic updates, HEPPO-GAE addresses GAE’s computational demands with a parallel, pipelined architecture implemented on a single System-on-Chip (SoC). This design allows for the adaptation of various hardware accelerators tailored for different PPO phases. A key innovation is our strategic standardization technique, which combines dynamic reward standardization and block standardization for values, followed by 8-bit uniform quantization. This method stabilizes learning, enhances performance, and manages memory bottlenecks, achieving a 4x reduction in memory usage and a 1.5x increase in cumulative rewards. We propose a solution on a single SoC device with programmable logic and embedded processors, delivering throughput orders of magnitude higher than traditional CPU-GPU systems. Our single-chip solution minimizes communication latency and throughput bottlenecks, significantly boosting PPO training efficiency. Experimental results show a 30% increase in PPO speed and a substantial reduction in memory access time, underscoring HEPPO-GAE’s potential for broad applicability in hardware-efficient reinforcement learning algorithms.
@@ I-C Related Work
This section discusses key contributions in hardware acceleration for RL, aligning with our work on HEPPO-GAE by addressing the computational demands of RL algorithms through innovative hardware designs, quantization techniques, and memory management strategies.
@@ II-C2 Quantization of Values
III HEPPO-GAE Architecture Details
@@ III-C System Architectue
Figure 5 (a) shows the micro-architecture of HEPPO-GAE, which consists of Rewards Loaders (ReLs), Values Loaders (VaLs), and compute Processing Elements (PEs) forming a one-dimensional systolic array with N rows.
Figure 5: HEPPO-GAE architecture consisting of Rewards Loaders (ReLs), Values Loaders (VaLs), PEs, system crossbar, and BRAM stack memory
Parallelization: Rows in the systolic array run concurrently and independently, each processing distinct vectors from different agents assigned by a round-robin fashion. When one row finishes, it gets a new set of vectors. This parallel architecture enhances HEPPO-GAE’s efficiency and scalability. While the BRAM stack memory enables substantial data transfers, a crossbar network ensures robust connections between ReLs, VaLs, and PEs to the BRAM stack memory.
@@ V-C Summary of Rewards and Values Quantization Approaches
V-D Hardware Implementation of HEPPO-GAE
A parameterized Verilog model of HEPPO-GAE’s proposed pipelined architecture, as described in Subsection III-C, has been developed with a data width of 32 bits (after fetching and de-quantizing the elements). The AMD-Xilinx Zynq® UltraScale+™ MPSoC ZCU106 Evaluation Kit was chosen to host our implementation. This device integrates a quad-core Arm® Cortex™-A53 processing system (PS) and a dual-core Arm Cortex-R5 real-time processor, providing the necessary computing for running the environment. The FPGA fabric within the Programmable Logic (PL) provides extensive resources for custom logic implementation, including neural network inference and GAE computation.
@@ VI Conclusion and Future Work
We introduced HEPPO-GAE (Hardware-Efficient Proximal Policy Optimization), an FPGA-based implementation designed to accelerate the GAE stage of the PPO algorithm. HEPPO-GAE utilizes dynamic reward standardization and 8-bit uniform quantization, reducing memory usage by 4x and increasing cumulative rewards by 50%.
@@
HEPPO-GAE leverages AMD-Xilinx Zynq UltraScale+ MPSoC’s capabilities, integrating environment simulation, neural network inference, backpropagation, and GAE computation on a single SoC. This minimizes communication latency and optimizes data handling which originally accounted for around 11% of the training.
Further incremental optimizations of HEPPO-GAE ’s SoC are possible. Overclocking techniques [18] and bit-serial computation [19] in FPGAs can be employed to accelerate overall processing. To mitigate power consumption, multiple clock domains can be implemented for the ARM cores, the DNN, and the GAE calculations. High-performance clock-domain crossing (CDC) FIFOs can facilitate faster data transfers [20, 21, 22, 23, 24]. Additionally, hardware-efficient data compression methods, optimized for deep learning workloads, can be leveraged to minimize data transfers [25, 26, 27].